# Joyce structures from quadratic differentials on the sphere

### v1

@@ 8. $\Omega$ as an $\mathcal{O}(2)$-valued twistor $2$-form
11-form $\Psi$ has poles only at $x\in\{\infty,w_{1},...,w_{N}\}$, corresponding to poles of $\phi$, and the pairs of poles corresponding to $x=q_{I}$ for I=1,,nI=1,...,n. Accordingly, $\Omega$ may be evaluated from a sum of residues at these points. Near the poles of $\phi$ we may expand
@@
$\Psi=\frac{\psi}{\hbar}+\sigma+\hbar\tau+O(\hbar^{2})$ (8.2)
@@
$\sigma$ $\coloneqq\frac{Q_{1}(x)}{2y_{0}}dx$ (8.3)
@@
The $1$-form $\sigma$ (previously introduced in §5) has constant residues at simple poles at the pairs of points corresponding to $q_{I}$ for I=1,,2nI=1,...,2n and no other poles. In particular, $U(\sigma):=U(Q_{1}(x)/2y_{0})dx$ represents a class in $H^{1}(\Sigma_{0}(p),\mathbb{C})$ for $U\in T_{\xi}X$. The $1$-form $\tau$ will have poles at the zeroes of $\psi$, with seemingly no useful condition on the residues, which makes it more difficult to handle.
@@
We will see that the upshot of Lemma 8.1 is that we can evaluate the contribution to $\Omega$ from residues at the poles of $\phi$ purely from the $1$-forms $\psi_{0},\sigma,\tau$ which are defined on $\Sigma_{0}(p)$. The same argument for the validity of the local expansion (8.2) does not work in the neighbourhood of a point corresponding to $x=q_{I}$ but the required terms in the Laurent expansion to evaluate the contribution of a residue here to $\Omega$ are
@@
$\Omega=\frac{\Omega_{-}}{\hbar^{2}}+\frac{i\Omega_{I}}{\hbar}+\Omega_{+}$ (8.7)
@@
$\Omega_{-}(U,V)$ $=2\pi i\sum_{\alpha}\underset{x=w_{\alpha}}{\mathrm{Res}}\big{(}d^{-1}U(\psi)V(\psi)\big{)}=\omega(U,V)$ (8.8)
@@
$={\langle U(\psi)|V(\sigma)\rangle_{0}+\langle U(\sigma)|V(\psi)\rangle_{0}}$ (8.10)
@@
$=$ $\underset{x=w_{\alpha}}{\text{Res}}\Bigg{(}d^{-1}U\Bigg{(}\frac{\psi}{\hbar}+{\sigma}\Bigg{)}V\Bigg{(}\frac{\psi}{\hbar}+{\sigma}\Bigg{)}+{d^{-1}U(\psi)V(\tau)+d^{-1}U(\tau)V(\psi)}\hskip-2.84526pt\Bigg{)}$ (8.14)
@@
$\underset{x=w_{\alpha}}{\mathrm{Res}}\big{(}d^{-1}U(\sigma)V(\sigma)\big{)}=0$ (8.15)
@@
Let $\Omega$ define a $2$-form on $X\times\mathbb{CP}^{1}$ by letting $\hbar=u^{1}/u^{0}$ where $u^{0}:u^{1}]\in\mathbb{CP}^{1}$. From $L(\hbar)\subseteq\ker\Omega$ (Proposition 6.2), and Proposition 8.2 above, we may (Proposition 2.5) interpret $\Omega$ as a closed $\mathcal{O}(2)$-valued relative $2$-form on the twistor space $Z=X\times\mathbb{CP}^{1}/L$.
@@
Let $L(\hbar)$ be the distribution given by the isomonodromic flows for fixed $\hbar$ and $L$ the associated twistor distribution. Then interpreting $\Omega$ as a $2$-form on $X\times\mathbb{CP}^{1}$ we have
$\operatorname{ker}\Omega=L\oplus p_{2}^{*}T\mathbb{CP}^{1}$ (8.17)
where $p_{2}$ is the projection $X\times\mathbb{CP}^{1}\to\mathbb{CP}^{1}$.
@@
As $\Omega$ defines a twistor $2$-form $\ker\Omega$ must be of rank at least $2n+1$. By Lemma 8.3 we see that $\ker\Omega$ is transverse to $VX$ for all $\hbar\in\mathbb{C}^{*}$ and therefore of rank at most $2n+1$. ∎
The above corollary implies that $\Omega$ has maximal rank as a relative $\mathcal{O}(2)$-valued $2$-form on $Z$. Altogether we have proven:
@@ 9. Symmetries of $g$ and the Joyce structure
$\hat{\iota}^{*}I=I,\quad\hat{\iota}^{*}(J\pm iK)=-(J\pm iK),\quad\hat{\iota}^{*}g=-g,$ (9.1)
@@
The first two equalities follow immediately from the definitions of the complex structures (2.3) and that, after inspecting the formulae for the isomonodromic flows in Appendix A, the vector fields $V_{I},V_{i}^{(\alpha)}$ are odd under simultaneously changing the signs of each $p_{J}$ while the $U_{I},U_{i}^{(\alpha)}$ remain unchanged.
For the last equality note that $\psi$ is invariant under $\tilde{\iota}$ since it depends only on the projection to $M$ while inspecting the formula for $Q_{1}$ we see $\sigma\mapsto-\sigma$. The expansion (8.6) shows that the $\hbar^{-1}$ term in the contribution to $\Omega$ from residues at points corresponding to $x=q_{I}$ is odd. Accordingly, we see that $\hat{\iota}^{*}\Omega_{I}=-\Omega_{I}$ which, using the definition $\Omega_{I}=g(I\cdot,\cdot)$ together with the first equality shows the last equality. ∎
@@
The vector field $E$ on $\tilde{X}$ given by
@@
satisfies
$\mathcal{L}_{E}I=0,\quad\mathcal{L}_{E}(J\pm iK)=\mp(J\pm iK),\quad\mathcal{L}_{E}g=g,$ (9.3)
@@
By inspecting the homogeneity of $Q_{0}(x),Q_{1}(x),Q_{2}(x)$ with respect to the parameters we see
@@
Integrating by parts yields
$[E(\Psi)]=\Bigg{[}\frac{1}{y}\Bigg{(}\frac{Q_{0}(x)}{\hbar^{2}}+\frac{Q_{1}(x)}{2\hbar}\Bigg{)}dx\Bigg{]}=:[\Phi]$ (9.6)
Treating $\hbar$ as a fixed constant and using Cartan’s formula this implies
$(\mathcal{L}_{E}\Omega)(U,V)=\langle U(\Phi)|V(\Psi)\rangle-\langle V(\Phi)|U(\Psi)\rangle.$ (9.7)
Near the poles of $Q_{0}(x)$ we expand in orders of $\hbar$
$\frac{1}{y}\Bigg{(}\frac{Q_{0}(x)}{\hbar^{2}}+\frac{Q_{1}(x)}{2\hbar}\Bigg{)}dx=\frac{\psi}{\hbar}-\hbar\tau+O(\hbar^{2})$ (9.8)
where the omitted terms have zeroes of order at least $2m_{\alpha}-2$ at $x=w_{\alpha}$. Taking a Laurent expansion near $x=q_{I}$ we have
@@
$=2\pi i\sum_{\alpha}\underset{x=w_{\alpha}}{\mathrm{Res}}\bigg{(}\hskip-2.84526ptd^{-1}U\bigg{(}\frac{\psi}{\hbar}-\hbar\tau\bigg{)}V\bigg{(}\frac{\psi}{\hbar}+\sigma+\hbar\tau\bigg{)}\hskip-4.2679pt\bigg{)}+2\pi i\sum_{I=1}^{n}\frac{dp_{I}(U)dq_{I}(V)}{\hbar}+\cdots$ (9.10)
where the omitted terms are proportional to $dq_{I}(U)dq_{I}(V)$ do not contribute to the result of anti-symmetrising the above in $U$ and $V$. Comparing to (8.8) and (8.9) we obtain
@@
This is an equality which holds on an open dense region of $X$ and hence on $X$. It follows
$\mathcal{L}_{E}\Omega_{-}=2\Omega_{-},\quad\mathcal{L}_{E}\Omega_{I}=\Omega_{I},\quad\mathcal{L}_{E}\Omega_{+}=0.$ (9.12)
These, together with the quaternion relations and the hermiticity of $g=-\Omega_{I}(I\cdot,\cdot)$ are equivalent to (9.3). ∎
@@
$\mathbb{T}$ is equipped with period coordinates $\{z_{a},\theta_{a}\}_{a=1}^{2n}$, where $\{z_{a}\}_{a=1}^{2n}$ are the period coordinates on $M$ and given $p\in M$ and $\omega\in H^{1}(\Sigma_{0}(p),\mathbb{C})\cong T_{p}M$
$\theta_{a}=\oint_{\gamma_{a}}\omega\ \,\mod 2\pi i.$ (9.13)
$\theta_{a}$ is the usual fibre (tangent) coordinate associated with $z_{a}$ in the sense that
$\theta^{a}\Bigg{(}\sum_{b=1}^{2n}V^{b}\frac{\partial}{\partial z^{b}}\Bigg{)}=\oint_{\gamma_{a}}\sum_{b=1}^{2n}V^{b}\frac{\partial y_{0}}{\partial z_{b}}dx=\sum_{b=1}^{2n}V^{b}\frac{\partial z^{a}}{\partial z^{b}}=V^{a}.$ (9.14)
@@
$\theta_{a}=\oint_{\gamma_{a}}\varpi$ (9.15)
@@
is the general one-form with simple poles with residue $1$ at the $n$ points $(q_{I},p_{I})$ residue n-n at $\infty$, and no other poles. Therefore, the periods $\theta_{a}$ are defined up to addition by $2\pi ik$ for $k\in\mathbb{Z}$. There are unique $t_{a}\in\mathbb{C}$ for a=1,,na=1,...,n such that
$2\sigma=\sum_{I=1}^{n}\nu_{I}+\sum_{a=1}^{n}t_{a}\omega_{a}.$ (9.18)
Here $\nu_{I}$ and $\omega_{a}$ are the normalised meromorphic $1$-forms as in [4]: The normalisation condition is that $\nu_{I}$ has simple poles with residues $\pm 1$ at points $(q_{I},p_{I})$ and $\infty$, respectively, and has $A$-cycles that are multiples of $2\pi i$. The $\omega_{a}$ are holomorphic, normalised so
@@
We have the following identities (from [4] Section 4.3):
$\oint_{A_{b}}\sum_{I=1}^{n}\nu_{I}$ $=0,$ (9.20)
@@
$\theta_{a}=\oint_{A_{a}}\varpi=2\pi it_{a}\,\ \operatorname{mod}2\pi i.$ (9.22)
@@
$\theta_{a+n}=\oint_{B_{a}}\varpi=\sum_{b=1}^{n}B_{ab}t_{b}+\sum_{I=1}^{n}\int_{\infty}^{(q_{I},p_{I})}\omega_{a}\ \,\operatorname{mod}2\pi i.$ (9.23)
where $B_{ab}$ is the matrix of $B$-periods for the basis of holomorphic $1$-forms $\{\omega_{a}\}_{a=1}^{n}$. Now the second term on the right-hand side is precisely the $a$-th components of the Abel-Jacobi map on the divisor depending on $\xi\in X_{p}$
@@
We now need some standard facts about the Abel-Jacobi map and the reduced representations of divisors. A comprehensive reference is [21] §2. The Abel-Jacobi map is a biholomorphism from the group $\operatorname{Pic}_{0}(\Sigma_{0}(p))$ of degree-zero divisors modulo principal divisors to the Jacobian variety $J(\Sigma_{0}(p))=\mathbb{C}^{n}/\Gamma$ where $\Gamma$ is the integer lattice in $\mathbb{C}^{n}$ generated by the $2n$-vectors $e_{a},\sum_{b}B_{ab}e_{b}\in\mathbb{C}^{n}$ from a=1,,na=1,...,n. The Abel-Jacobi map is surjective when restricted to divisors of the form $\sum_{I=1}^{n}(r_{I}-\infty)$ for points $r_{I}\in\Sigma_{0}(p)$ not necessarily distinct. Recall the conditions on the points $(q_{I},p_{I})$ was that the $q_{I}$ were distinct and away from the branch points. This gives an open dense subset of $\text{Pic}_{0}(\Sigma_{0}(p))$. Accordingly, $\Theta$ has an image an open dense subset of a torus bundle $\mathbb{T}\to M$. Now the divisors (9.24) are sent to different points on the Jacobian, up to the action of the symmetric group on the $q_{I}$, since otherwise the difference of two such divisors would represent a principal divisor of a meromorphic function with less than or equal to $g$ poles (counting multiplicity) away from $\infty$ and with a set of poles none of which are mapped to each other by the covering involution $y\mapsto-y$. It follows that $\Theta$ is an immersion. ∎
@@
The map $\tilde{\iota}:\tilde{X}\to\tilde{X}$ induces $\iota:\mathbb{T}\to\mathbb{T}$ given by
@@
$\theta_{a}\circ\tilde{\iota}=\pi ik-\oint_{\gamma_{a}}\sigma=-\bigg{(}\pi ik+\oint_{\gamma_{a}}\sigma\bigg{)}+2\pi ik=-\theta_{a}+2\pi ik$ (9.27)
@@
In the period coordinates $\{z_{a},\theta_{a}\}_{a=1}^{2n}$ the push-forward of $E$ is given by
$E=\sum_{a=1}^{2n}z_{a}\frac{\partial}{\partial z_{a}}.$ (9.28)
@@
$E(Q_{0}(x))+\frac{2}{2m_{\infty}-3}x\frac{\partial}{\partial x}Q_{0}(x)=\frac{4m_{\infty}-10}{2m_{\infty}-3}Q_{0}(x)$ (9.29)
@@
$E(y)dx=\frac{2m_{\infty}-5}{2m_{\infty}-3}\,ydx-\frac{1}{2m_{\infty}-3}\frac{xQ^{\prime}_{0}(x)}{y}\,dx.$ (9.31)
Integrating the last term by parts yields $W(\psi)]=[\psi]\in H^{1}(\Sigma_{0}(p),\mathbb{C})$. In particular,
$E(z_{a})=\int_{\gamma_{a}}E(\psi)=z_{a}$ (9.32)

### v2

@@ 8. $\Omega$ as an $\mathcal{O}(2)$-valued twistor $2$-form
The $1$-form $\Psi$ has poles only at $x\in\{\infty,w_{1},...,w_{N}\}$, corresponding to poles of $\phi$, and the pairs of poles corresponding to $x=q_{I}$ for I=1,,nI=1,...,n. Accordingly, $\Omega$ may be evaluated from a sum of residues at these points. Near the poles of $\phi$ we may expand
@@
$\Psi=\frac{\psi}{\hbar}+\sigma+\hbar\tau+O(\hbar^{2}),$ (8.2)
@@
$\sigma$ $\coloneqq\frac{Q_{1}(x)}{2y_{0}}dx,$ (8.3)
@@
The $1$-form $\sigma$ (previously introduced in §5) has constant residues at simple poles at the pairs of points corresponding to $q_{I}$ for I=1,,2nI=1,...,2n and no other poles. In particular, $U(\sigma):=U(Q_{1}(x)/2y_{0})dx$ represents a class in $H^{1}(\Sigma_{0}(p),\mathbb{C})$ for $U\in T_{\xi}X$.
@@
We will see that the upshot of Lemma 8.1 is that we can evaluate the contribution to $\Omega$ from residues at the poles of $\phi$ purely from the $1$-forms $\psi,\sigma,\tau$ which are defined on $\Sigma_{0}(p)$. The same argument for the validity of the local expansion (8.2) does not work in the neighbourhood of a point corresponding to $x=q_{I}$ but the required terms in the Laurent expansion to evaluate the contribution of a residue here to $\Omega$ are (see also (6.16))
@@
$\Omega=\frac{\Omega_{-}}{\hbar^{2}}+\frac{i\Omega_{I}}{\hbar}+\Omega_{+},$ (8.7)
@@
$\Omega_{-}(U,V)$ $=2\pi i\sum_{\alpha}\underset{x=w_{\alpha}}{\mathrm{Res}}\big(d^{-1}U(\psi)V(\psi)\big)=\omega(U,V),$ (8.8)
@@
$={\langle U(\psi)|V(\sigma)\rangle_{0}+\langle U(\sigma)|V(\psi)\rangle_{0}},$ (8.10)
@@
$=$ $\underset{x=w_{\alpha}}{\text{Res}}\Bigg(d^{-1}U\Bigg(\frac{\psi}{\hbar}+{\sigma}\Bigg)V\Bigg(\frac{\psi}{\hbar}+{\sigma}\Bigg)+{d^{-1}U(\psi)V(\tau)+d^{-1}U(\tau)V(\psi)}\hskip-2.84526pt\Bigg),$ (8.14)
@@
$\underset{x=w_{\alpha}}{\mathrm{Res}}\big(d^{-1}U(\sigma)V(\sigma)\big)=0,$ (8.15)
@@
Let $\Omega$ define a $2$-form on $X\times\mathbb{CP}^{1}$ by letting $\hbar=u^{1}/u^{0}$ where $u^{0}:u^{1}]\in\mathbb{CP}^{1}$. From $L(\hbar)\subseteq\ker\Omega$ (Proposition 6.2) and Proposition 8.2 above, we may (Proposition 2.5) interpret $\Omega$ as a closed $\mathcal{O}(2)$-valued relative $2$-form on the twistor space $Z=X\times\mathbb{CP}^{1}/L$.
@@
Interpreting $\Omega$ as a $2$-form on $X\times\mathbb{CP}^{1}$ we have
$\operatorname{ker}\Omega=L\oplus p_{2}^{*}T\mathbb{CP}^{1},$ (8.17)
where $L$ is the twistor distribution and $p_{2}$ is the projection $X\times\mathbb{CP}^{1}\to\mathbb{CP}^{1}$.
@@
Lemma 6.2 states $L\subseteq\operatorname{ker}\Omega$ and Proposition 7.3 implies $L$ is of rank $2n$. Clearly $p_{2}^{*}T\mathbb{CP}^{1}\subseteq\ker\Omega$. By Lemma 8.3, we see that $\ker\Omega$ is transverse to $VX$ for all $\hbar\in\mathbb{C}^{*}$ and therefore of rank at most $2n+1$. ∎
The above corollary implies that $\Omega$ has maximal rank as a relative $\mathcal{O}(2)$-valued $2$-form on $Z$. Applying Theorem 2.6, we have proven:
@@ 9. Symmetries of $g$ and the Joyce structure
$\hat{\iota}^{*}I=I,\quad\hat{\iota}^{*}(J\pm iK)=-(J\pm iK),\quad\hat{\iota}^{*}g=-g.$ (9.1)
@@
The first two equalities follow immediately from the definitions of the complex structures (2.3) and that, after inspecting the formulae for the isomonodromic flows in Appendix B, the vector fields $V_{I},V_{i}^{(\alpha)}$ are odd under simultaneously changing the signs of each $p_{J}$ while the $U_{I},U_{i}^{(\alpha)}$ remain unchanged.
For the last equality note that $\psi$ is invariant under $\hat{\iota}$ since it depends only on the projection to $M$ while inspecting the formula for $Q_{1}$ we see $\sigma\mapsto-\sigma$. The expansion (8.6) shows that the $\hbar^{-1}$ term in the contribution to $\Omega$ from residues at points corresponding to $x=q_{I}$ is odd. Accordingly, we see that $\hat{\iota}^{*}\Omega_{I}=-\Omega_{I}$ which, using the definition $\Omega_{I}=g(I\cdot,\cdot)$ together with the first equality shows the last equality. ∎
To demonstrate the hyper-Kähler structure has the appropriate homogeneity properties, the following will be helpful:
Lemma 9.2.
A complex hyper-Kähler structure admits a vector field $E$ satisfying
$\mathcal{L}_{E}\Omega_{-}=2\Omega_{-},\quad\mathcal{L}_{E}\Omega_{I}=\Omega_{I},\quad\mathcal{L}_{E}\Omega_{+}=0,$ (9.2)
if and only if
$\mathcal{L}_{E}I=0,\quad\mathcal{L}_{E}(J\pm iK)=\mp(J\pm iK),\quad\mathcal{L}_{E}g=g.$ (9.3)
Proof.
That (9.3) implies (9.2) is obvious from the Leibniz rule and the definitions of the $2$-forms. Conversely, suppose (9.2) holds. Using the Leibniz rule and the definitions of the relevant $2$-forms we have equations:
$g(\mathcal{L}_{E}I\cdot,\cdot)+(\mathcal{L}_{E}g)(I\cdot,\cdot)$ $=g(I\cdot,\cdot),$ (9.4)
$g(\mathcal{L}_{E}J\cdot,\cdot)+(\mathcal{L}_{E}g)(J\cdot,\cdot)-i(g(\mathcal{L}_{E}K\cdot,\cdot)+(\mathcal{L}_{E}g)(K\cdot,\cdot))$ $=2(g(J\cdot,\cdot)-ig(K\cdot,\cdot)),$ (9.5)
$g(\mathcal{L}_{E}J\cdot,\cdot)+(\mathcal{L}_{E}g)(J\cdot,\cdot)+i(g(\mathcal{L}_{E}K\cdot,\cdot)+(\mathcal{L}_{E}g)(K\cdot,\cdot))$ $=0.$ (9.6)
Rewrite the last two equations as
$g(\mathcal{L}_{E}J\cdot,\cdot)+(\mathcal{L}_{E}g)(J\cdot,\cdot)$ $=g(J\cdot,\cdot)-ig(K\cdot,\cdot),$ (9.7)
$g(\mathcal{L}_{E}K\cdot,\cdot)+(\mathcal{L}_{E}g)(K\cdot,\cdot)$ $=i(g(J\cdot,\cdot)-ig(K\cdot,\cdot)).$ (9.8)
Precompose the first and second equations with $K$ and $J$ respectively, then add and subtract the two equations to obtain
$g((\mathcal{L}_{E}J)K\cdot,\cdot)+g((\mathcal{L}_{E}K)J\cdot,\cdot)$ $=0,$ (9.9)
$g((\mathcal{L}_{E}J)K\cdot,\cdot)-g((\mathcal{L}_{E}K)J\cdot,\cdot)$ $=-2(\mathcal{L}_{E}g)(I\cdot,\cdot)+2g(I\cdot,\cdot)+2ig(\cdot,\cdot).$ (9.10)
Putting the above together with (9.4) and $\mathcal{L}_{E}I=(\mathcal{L}_{E}J)K+J(\mathcal{L}_{E}K)$, we see
$2ig(\cdot,\cdot)+2g(J(\mathcal{L}_{E}K),\cdot)=0$ (9.11)
so that $\mathcal{L}_{E}K=iJ$. Substituting this into (9.8) gives
$(\mathcal{L}_{E}g)(K\cdot,\cdot)=g(K\cdot,\cdot)$ (9.12)
and hence, after precomposing with $K$, we obtain $\mathcal{L}_{E}g=g$. Using this, it is straightforward, again using the Leibniz rule, to verify the rest of (9.3). ∎
@@
The vector field $E$ on $X$ given by
@@
satisfies (9.3).
@@
By inspecting the homogeneity of $Q_{0}(x),Q_{1}(x),Q_{2}(x)$ with respect to the parameters, we see
@@
Integrating by parts using $d(xy)=ydx+xQ^{\prime}(x)/2y$ yields
$[E(\Psi)]=\Bigg[\frac{1}{y}\Bigg(\frac{Q_{0}(x)}{\hbar^{2}}+\frac{Q_{1}(x)}{2\hbar}\Bigg)dx\Bigg]=:[\Phi].$ (9.16)
Treat $\hbar$ as a fixed constant and use Cartan’s formula:
$(\mathcal{L}_{E}\Omega)(U,V)$ $=d(\Omega(E,\cdot))(U,V)$
$=U(\Omega(E,V))-V(\Omega(E,U))-\Omega(E,[U,V])$
$=U(\langle\Phi|V(\Psi)\rangle)-V(\langle\Phi|U(\Psi)\rangle)-\langle\Phi|[U,V](\Psi)\rangle$
$=\langle U(\Phi)|V(\Psi)\rangle-\langle V(\Phi)|U(\Psi)\rangle.$ (9.17)
Note that
$\Phi=\hbar\frac{\partial\Psi}{\partial\hbar}dx.$ (9.18)
We combine this observation with (8.2). Near the poles of $Q_{0}(x)$, expand in orders of $\hbar$
$\frac{1}{y}\Bigg(\frac{Q_{0}(x)}{\hbar^{2}}+\frac{Q_{1}(x)}{2\hbar}\Bigg)dx=\frac{\psi}{\hbar}-\hbar\tau+O(\hbar^{2}),$ (9.19)
where the omitted terms have (see Lemma (8.1)) zeroes of order at least $2m_{\alpha}-2$ at $x=w_{\alpha}$. Similarly, from (8.6) we have the Laurent expansion near $x=q_{I}$:
@@
$=2\pi i\sum_{\alpha}\underset{x=w_{\alpha}}{\mathrm{Res}}\bigg(\hskip-2.84526ptd^{-1}U\bigg(\frac{\psi}{\hbar}-\hbar\tau\hskip-2.84526pt\bigg)V\bigg(\frac{\psi}{\hbar}+\sigma+\hbar\tau\hskip-2.84526pt\bigg)\hskip-4.2679pt\bigg)+2\pi i\sum_{I=1}^{n}\frac{dp_{I}(U)dq_{I}(V)}{\hbar}+\cdots\hskip-1.42262pt,$ (9.21)
where the omitted terms are proportional to $dq_{I}(U)dq_{I}(V)$ and do not contribute to the result of anti-symmetrising the above in $U$ and $V$. Comparing to (8.8) and (8.9) we obtain the following equality, which holds on an open dense region of $X$ and hence on $X$:
@@
That this holds to all orders in $\hbar$ is equivalent to (9.2), which implies (9.3). ∎
@@
$\mathbb{T}$ is equipped with period coordinates $\{z^{a},\theta^{a}\}_{a=1}^{2n}$, where $\{z^{a}\}_{a=1}^{2n}$ are the period coordinates on $M$ and given $p\in M$ and $\omega\in H^{1}(\Sigma_{0}(p),\mathbb{C})\cong T_{p}M$
$\theta^{a}=\oint_{\gamma_{a}}\omega\ \,\mod 2\pi i.$ (9.23)
$\theta^{a}$ is the usual fibre (tangent) coordinate associated with $z^{a}$ in the sense that
$\theta^{a}\Bigg(\sum_{b=1}^{2n}V^{b}\frac{\partial}{\partial z^{b}}\Bigg)=\oint_{\gamma_{a}}\sum_{b=1}^{2n}V^{b}\frac{\partial y_{0}}{\partial z^{b}}dx=\sum_{b=1}^{2n}V^{b}\frac{\partial z^{a}}{\partial z^{b}}=V^{a}.$ (9.24)
@@
$\theta^{a}=-\oint_{\gamma_{a}}\varpi$ (9.25)
@@
is the general one-form with simple poles with residue $1$ at the $n$ points $(q_{I},p_{I})$, residue n-n at $\infty$, and no other poles. Therefore, the functions $\theta^{a}$ are defined up to addition by $2\pi ik$ for $k\in\mathbb{Z}$. There are unique $t_{a}\in\mathbb{C}$ for a=1,,na=1,...,n such that
$\varpi=\sum_{I=1}^{n}\nu_{I}+\sum_{a=1}^{n}t_{a}\omega_{a}.$ (9.28)
Here $\nu_{I}$ and $\omega_{a}$ are normalised meromorphic $1$-forms as in [5]: The normalisation condition is that $\nu_{I}$ has simple poles with residues $\pm 1$ at points $(q_{I},p_{I})$ and $\infty$, respectively, and we have chosen a set of representative $A$-cycles for which the periods vanish. The $\omega_{a}$ are holomorphic, normalised so
@@
We have the following identities ([5] Section 4.3), for any set of representative cycles:
$\oint_{A_{b}}\sum_{I=1}^{n}\nu_{I}$ $=0\ \,\operatorname{mod}2\pi i,$ (9.30)
@@
$-\theta^{a}=\oint_{A_{a}}\varpi=2\pi it_{a}\,\ \operatorname{mod}2\pi i.$ (9.32)
@@
$-\theta^{a+n}=\oint_{B_{a}}\varpi=\sum_{b=1}^{n}B_{ab}t_{b}+\sum_{I=1}^{n}\int_{\infty}^{(q_{I},p_{I})}\omega_{a}\ \,\operatorname{mod}2\pi i$ (9.33)
where $B_{ab}$ is the matrix of $B$-periods for the basis of holomorphic $1$-forms $\{\omega_{a}\}_{a=1}^{n}$. Now, the second term on the right-hand side is precisely the $a$th components of the Abel-Jacobi map on the divisor depending on $\xi\in X_{p}$
@@
We now need some standard facts about the Abel-Jacobi map and the reduced representations of divisors. Comprehensive references are [1, 24]. The Abel-Jacobi map is a biholomorphism from the group $\operatorname{Pic}_{0}(\Sigma_{0}(p))$ of degree-zero divisors modulo principal divisors to the Jacobian variety $J(\Sigma_{0}(p))=\mathbb{C}^{n}/\Gamma$ where $\Gamma$ is the integer lattice in $\mathbb{C}^{n}$ generated by the $2n$-vectors $e_{a},\sum_{b}B_{ab}e_{b}\in\mathbb{C}^{n}$ from a=1,,na=1,...,n. The Abel-Jacobi map is surjective when restricted to classes represented by divisors $\sum_{I=1}^{n}(r_{I}-\infty)$ for points $r_{I}\in\Sigma_{0}(p)$ not necessarily distinct. Recall the conditions on the points $(q_{I},p_{I})$ was that the $q_{I}$ were distinct and away from the branch points. This gives an open dense subset of $\text{Pic}_{0}(\Sigma_{0}(p))$. Accordingly, $\Theta$ has image an open dense subset of a torus bundle $\mathbb{T}\to M$. The differential of the Abel-Jacobi map has corank that is the dimension of the space of holomorphic $1$-forms vanishing at the points $(q_{I},p_{I})$, which may be calculated as, using the Riemann-Roch theorem, $h^{0}(\Sigma_{0}(p),\mathcal{O}(\tilde{D}(\xi)))-1$ where $\tilde{D}(\xi)=\sum_{I=1}^{n}(q_{I},p_{I})$. Now we use the standard fact that a divisor of this form consisting of distinct points is special if and only if it contains points interchanged by the covering involution. So $\tilde{D}(\xi)$ is not special and $h^{0}(\Sigma_{0}(p),\mathcal{O}(\tilde{D}(\xi)))=1$. It follows that $\Theta$ is an immersion. ∎
@@
The map $\hat{\iota}:X\to X$ induces $\iota:\mathbb{T}\to\mathbb{T}$ given by
@@
$\theta^{a}\circ\hat{\iota}=\pi ik+\oint_{\gamma_{a}}\sigma=-\bigg(\pi ik-\oint_{\gamma_{a}}\sigma\bigg)+2\pi ik=-\theta^{a}+2\pi ik$ (9.37)
@@
In the period coordinates $\{z^{a},\theta^{a}\}_{a=1}^{2n}$ the push-forward of $E$ is given by
$E=\sum_{a=1}^{2n}z^{a}\frac{\partial}{\partial z^{a}}.$ (9.38)
@@
$E(Q_{0}(x))+\frac{2}{2m_{\infty}-3}x\frac{\partial}{\partial x}Q_{0}(x)=\frac{4m_{\infty}-10}{2m_{\infty}-3}Q_{0}(x),$ (9.39)
@@
$E(y_{0})dx=\frac{2m_{\infty}-5}{2m_{\infty}-3}\,y_{0}dx-\frac{1}{2m_{\infty}-3}\frac{xQ^{\prime}_{0}(x)}{y_{0}}\,dx.$ (9.41)
Integrating the last term by parts yields $E(\psi)]=[\psi]\in H^{1}(\Sigma_{0}(p),\mathbb{C})$. In particular,
$E(z^{a})=\int_{\gamma_{a}}E(\psi)=z^{a}$ (9.42)